# High Fidelity Text-Guided Music Generation and Editing via Single-Stage Flow Matching

### v1

@@ 3 Experimental setup
Model
MelodyFlow is a diffusion transformer of sizes 400M (small) and 1B (medium) parameters with U-shaped skip connections Bao et al. (2023). The model is conditioned via cross attention on a T5 representation (Raffel et al., 2020) computed from the text description of the music. The flow step is injected following (Hatamizadeh et al., 2023). Minibatch coupling is computed with torch-linear-assignement [^1] https://github.com/ivan-chai/torch-linear-assignment. MelodyFlow-small (resp. MelodyFlow-medium) is trained on latent representation sequences of 32 kHz mono (resp. 48 kHz stereo) segments of 10 (resp. 30) seconds, encoded at 20 Hz frame rate (resp. 25 Hz). From the codec perspective the only difference between encoding mono or stereo waveform is the number of input (resp. output) channels for the first (resp. last) convolution of the encoder (resp. decoder): 1 for mono and 2 for stereo. More details regarding audio representation and FM model training are provided in appendix A.1.
Inference and editing
For text-to-music inference we use the midpoint ODE solver from torchdiffeq with a step size of $0.03125$. Classifier free guidance of $5.0$ is chosen after grid search. For music editing we run our regularized inversion method with the euler ODE solver until $T_{edit}=0.04$ with the same classifier free guidance of $5.0$ applied in both ODE directions. Trajectory inversion estimation is run with $S=25,K=4,w_{k}=k-1$ and $\lambda_{KL}=0.2$.
@@ 3.1 Datasets
For the main text-to-music generation results we evaluate MelodyFlow and prior work on the MusicCaps dataset (Agostinelli et al., 2023). We compute objective metrics for MelodyFlow and report those from previous literature. Subjective evaluations are conducted on a subset of 198 examples from the genre-balanced set. For ablations we rely on an in-domain held out evaluation set different from that of (Copet et al., 2024), made of 8377 tracks. The same in-domain tracks are used for objective editing evaluations. Subjective evaluations of edits are run on a subset of 181 higher fidelity samples from our in-domain test set with LLM-assisted designed prompts (more details in appendix A.1.3).
Metrics
We evaluate MelodyFlow using both objective and subjective metrics, following the evaluation protocol of (Kreuk et al., 2022; Copet et al., 2024) for generation. Reported objective metrics are the Fréchet Audio Distance (FAD) (Roblek et al., 2019) with VGGish embeddings (Hershey et al., 2017), the Kullback–Leibler divergence (KLD) with PASST audio encoder (Koutini et al., 2021) and CLAP [^2] https://github.com/LAION-AI/CLAP cosine similarity (Elizalde et al., 2023). For music editing evaluations we compute the average L2 distance between the original and edited latent sequences (LPAPS (Iashin and Rahtu, 2021)), FADedit between the distribution of source and edited samples and CLAPedit between the edited audio and the editing prompt. Subjective evaluations relate to (i) overall quality (OVL), and (ii) relevance to the text input (REL), both using a Likert scale (from 1 to 5). Additionally for music editing evaluations we report (iii) editing consistency (CON) where raters are asked to score the level of consistency between the editing and original samples, considering the provided editing prompt. Raters were recruited using the Amazon Mechanical Turk platform and all samples were normalized to -14dB LUFS (Series, 2011). For stereo samples objective evaluation the signal is down mixed into mono prior to metrics computation. For subjective ratings we keep the original audio format generated by each model.
@@ 4 Results
We first evaluate the model performance on text-to-music generation before investigating text-guided music editing. For text-to-music qualitative evaluations we compare MelodyFlow to three baselines: MusicGen, AudioLDM 2 and Stable-Audio. For MusicGen, AudioLDM 2 we use the available open source implementations and for Stable-Audio we use the public API (as of Wed. May 14 2024, AudioSpark 2.0 model version). For music editing qualitative evaluations we compare MelodyFlow to MusicGen-melody (specifically trained with chroma conditioning for the task of text-conditioned music editing), AudioLDM 2 with DDPM inversion (following (Manor and Michaeli, 2024)). Although Stable-Audio API offers a music editing setting, for practical reasons we were not able to submit hundreds of samples hence exclude it from our baselines.
4.1 Text-to-music generation
Text-to-music generation performance is reported in table 1. MelodyFlow achieves comparable performance with MusicGen, both lagging slightly behind Stable-Audio in terms of human preference. We do not report objective metrics on Stable-Audio as none were reported on the full MusicCaps benchmark Evans et al. (2024a). MelodyFlow achieves remarkable efficiency with only 64 inference steps.
[…]
Reference - - - 3.67$\pm$0.10 4.04$\pm$0.10 -
AudioLDM 2 3.1 $1.20$ $0.31$ 2.79$\pm$0.08 3.40$\pm$0.08 208
MusicGen-small 3.1 $1.29$ $0.31$ 1500
MusicGen-medium 3.4 $1.23$ $0.32$ 3.40$\pm$0.08 3.79$\pm$0.07 1500
Stable-Audio - - - 3.67$\pm$0.08 3.89$\pm$0.07 100
MelodyFlow-small 2.8 1.27 0.33 3.27$\pm$0.08 3.83$\pm$0.08 64
MelodyFlow-medium 3.5 1.30 0.31 3.41$\pm$0.08 3.77$\pm$0.07 64
@@ 4.2 Text-guided music editing
The table 2 presents the main music editing results. We compare MelodyFlow-medium with our inversion method to AudioLDM 2 with DDPM inversion and MusicGen-melody. MelodyFlow outperforms both baselines on the quality and text-fidelity axes. Consistency-wise MelodyFlow lags slightly behind AudioLDM 2. Indeed during our listening tests we observe that AudioLDM 2 with DDPM inversion sometimes only generates a distorted version of the original track. Averaging on the three axes MelodyFlow sets a new baseline for zero-shot music editing at test-time.
@@
4.3 Ablations
4.3.1 Codec bottleneck and framerate
We ablate on the bottleneck choice for a fixed frame rate of 50Hz by comparing RVQ ((Copet et al., 2024) setting), VAE and identity in Table 3. From the codec perspective our results indicate optimal reconstruction performance with the identity or VAE bottlenecks followed by RVQ. The same ranking applies for SI-SDR (Le Roux et al., 2019). Regarding FM model performance we report a decreasing in-domain FAD in the favor the VAE. Ablating on the VAE codec frame rate shows that 5 Hz is comparable to 50 Hz codec with RVQ regarding codec performance. FM model performance is closer to that of 50 Hz codec with identity bottleneck. For MelodyFlow-small we chose to use 20 Hz.
@@ 4.3.1 Codec bottleneck and framerate
4.3.2 Scaling for high-fidelity and duration versatility.
The table 4 reports the impact of moving from mono to stereo with the same MelodyFlow-medium model size. We also report the effect of applying a L-shaped attention mask during model training to support durations shorter than 30 seconds during inference. For each sequence of length L, we randomly select a segment boundary within the range $[0,L]$. Positions before the boundary can only attend to themselves in the self-attention, while positions after it attend to the entire sequence. The in-domain FAD is reported for 10s and 30s generated segments. Our results indicate that the L-mask helps supporting versatile duration with no penalty on full-length segments, unlocking faster inference for shorter segments. However moving from mono to stereo only slightly affects the generative model performance, likely due to our mono-based evaluation.
Table 4: Ablation on L-mask and stereo for MelodyFlow-large. Each variant is trained on 30s audio segments encoded with a 25 Hz frame rate codec trained on 48 kHz audio.
[…]
2 0.40 12.48 0.59 0.65
1.48 0.65
1 0.39 13.34 0.49 0.59
4.3.3 Flow matching design choices
Table 5 presents the impact of the key design choices in the flow matching model training. Considering our baseline implementation was initially designed for text-to-speech (Le et al., 2024), it included an infilling objective that was meant to help the model handle variable length sequences that are inherent to the speech domain. We report the in-domain test FAD and the last validation MSEloss computed from the EMA checkpoint. No loss value is report for the baseline as the infilling objective facilitates the task. For the second line of the table the reported loss is weighted by the logit-normal probability density function at each sampled flow step to be comparable with logit-normal sampling. With all methods combined the in-domain FAD is reduced to 0.39 from 0.53 and consistent with the loss decrease, which validates our design choices.
Table 5: Flow matching model design ablation. FAD (resp. MSE) is reported on the in-domain test (resp. validation) set. Baseline is adapted from (Le et al., 2024) but retrained on our music latents.
[…]
baseline 16 24 uniform .53 -
-- infilling 16 24 uniform .50 .8596
$+$ sampling 16 24 logit-normal .44 .8484
$+$ batch coupling 16 24 logit-normal .42 .8322
$+$ wider model 18 18 logit-normal .39 .8310
(a) Editing FADedit vs. $T_{edit}$.
(b) Editing CLAPedit vs. $T_{edit}$.
(c) Editing LPAPS vs. $T_{edit}$.
4.3.4 Flow matching inversion
In the Figures 1(a), 1(b) and 1(c) we report music editing objective metrics as a function of $T_{edit}$, comparing naive DDIM inversion with MelodyFlow. For MelodyFlow inversion we ablate on the original text description used in the backward ODE process ($c_{T_{edit}\longleftarrow{}1}\in\{\varnothing,c_{orig}\}$). DDIM with text conditioning is not shown due to its instability.We observe that our proposed regularized inversion outperforms DDIM inversion, irrespective of the use of the text conditioning in the inversion process. Using $c_{orig}$ to condition the inversion shows better text adherence (higher CLAPedit) at the detriment of quality (higher FADedit). Consistency with the original song is generally much higher (lower LPAPS) than when using DDIM inversion, which correlates with our listening tests. The S-shaped FAD curves of the Figure 1(a) indicate an inversion target optimum around $T_{edit}=0.06$.
@@ 4.3.4 Flow matching inversion
4.3.5 Music editing efficiency
We compare DDIM with MelodyFlow inversion using $T_{edit}=0$ and several efficiency budgets. FADedit (Figure 2(a)), CLAPedit (Figure 2(b)) and LPAPS (Figure 2(c)) are plotted as a function of NFE. A good trade off between speed, quality, text-adherence and consistency is achieved at 128 NFEs.
@@ 5 Related work
Audio representation
Recent advancements in neural codecs have seen the application of VQ-VAE on raw waveforms, incorporating a RVQ bottleneck as demonstrated in Zeghidour et al. (2021); Défossez et al. (2022), later refined as per Kumar et al. (2024). (Evans et al., 2024a) proposed a modification to this approach by replacing the RVQ with a VAE bottleneck to enhance the modeling of continuous representations. In addition, several recent audio generative models have adopted Mel-Spectrogram latent representations, coupled with a vocoder for reconstruction, as shown in the works of (Ghosal et al., 2023; Liu et al., 2023; Le et al., 2024). Furthermore, Défossez et al. (2022) introduced an additional layer of complexity by incorporating quantization to support discrete representation on top of the continuous representation.
Text-to-music generation
Models that operate on discrete representation are presented in the works of (Agostinelli et al., 2023; Copet et al., 2024; Ziv et al., 2023). Agostinelli et al. (2023) proposed a representation of music using multiple streams of tokens, which are modeled by a cascade of transformer decoders conditioned on a joint textual-music representation (Huang et al., 2022b). Copet et al. (2024) introduced a single-stage language model that operates on streams of discrete audio representations, supporting both 32 kHz mono and stereo. Ziv et al. (2023) replaced the language model with a masked generative single-stage non-autoregressive transformer. Schneider et al. (2023); Huang et al. (2023); Liu et al. (2023) use diffusion models. Schneider et al. (2023) utilized diffusion for both the generation model and the audio representation auto-encoder. Liu et al. (2023) trained a foundational audio generation model that supports music with latent diffusion, conditioned on autoregressively generated AudioMAE features (Huang et al., 2022a). Evans et al. (2024a, b) proposed an efficient long-form stereo audio generation model based on the latent diffusion of VAE latent representations. This model introduced timing embeddings conditioning as a method to better control the content and length of the generated music.
Music editing
Lin et al. (2024) proposed a parameter-efficient fine-tuning method for autoregressive language models to support music inpainting tasks. Garcia et al. (2023) developed a masked acoustic modeling approach for music inpainting, outpainting, continuation and vamping. Wu et al. (2023) fine-tuned a diffusion-based music generation model with melody, dynamics and rhythm conditioning. Novack et al. (2024) is a fine-tuning free framework for controlling pre-trained text-to-music diffusion models at inference-time via initial noise latent optimization. Zhang et al. (2024) investigated zero-shot text-guided music editing with conditional latent space and cross attention maps manipulation. Manor and Michaeli (2024) employs DDPM inversion (Huberman-Spiegelglas et al., 2023) for zero-shot unsupervised and text-guided audio editing.

### v2

@@ 3 Experimental setup
3.1 Model
MelodyFlow uses a DiT of sizes 400M (small) and 1B (medium) parameters with U-shaped skip connections Bao et al. (2023). The model is conditioned via cross attention on a T5 representation (Raffel et al., 2020) computed from the text description of the music. The model integrates a specific L-shaped self-attention mask meant to better generalize to different segment lengths during inference (appendix A.2.2). The flow step is injected following Hatamizadeh et al. (2023). Minibatch coupling is computed with torch-linear-assignement [^1] https://github.com/ivan-chai/torch-linear-assignment. MelodyFlow-small (resp. MelodyFlow-medium) is trained on latent representation sequences of 32 kHz mono (resp. 48 kHz stereo) segments of 10 (resp. 30) seconds, encoded at 20 Hz frame rate (resp. 25 Hz). From the codec perspective the only difference between encoding mono or stereo waveform is the number of input (resp. output) channels for the first (resp. last) convolution of the encoder (resp. decoder): 1 for mono and 2 for stereo. The appendix A.2.1 specifically investigates the impact of encoding stereo instead of mono signals on both reconstruction and generation performance. More details regarding audio representation and FM model implementation and training are provided in the appendix A.1.
3.2 Generation and editing
For text-to-music generation we use the midpoint ODE solver from torchdiffeq with a step size of $0.03125$. A classifier free guidance (CFG) of $4.0$ is chosen after grid search (appendix A.2.4). For music editing we use the same configuration for DDIM inversion. For ReNoise and MelodyFlow we use a longer step size of 0.04 to account for the additional forward passes induced by the reversible trajectory estimation. For DDIM inversion this gives a total of 64 inversion and 64 generation steps (e.g. forward passes through the DiT). For ReNoise and MelodyFlow the inversion takes 25 steps (each of them requires 4 iterations for the reversibility estimation) and 25 forward steps, for a total of 125. In summary MelodyFlow’s inversion is run with $S=25,K=4,w_{0}=w_{1}=0,w_{2}=2,w_{3}=3$ and $\lambda_{KL}=0.2$.
@@ 3.3 Datasets
For the main text-to-music generation results we evaluate MelodyFlow and prior work on the MusicCaps dataset (Agostinelli et al., 2023). We compute objective metrics for MelodyFlow and report those from previous literature. Subjective evaluations are conducted on a subset of 198 examples from the genre-balanced set. For ablations we rely on an in-domain held out evaluation set different from that of Copet et al. (2024), made of 8377 tracks. The same in-domain tracks are used for objective editing evaluations. Subjective evaluations of edits are run on a subset of 181 higher fidelity samples from our in-domain test set with LLM-assisted designed prompts (more details in appendix A.1.3).
3.4 Metrics
We evaluate MelodyFlow using both objective and subjective metrics following the evaluation protocol of Kreuk et al. (2022) and Copet et al. (2024) for generation. Reported objective metrics are the Fréchet Audio Distance (FAD) (Roblek et al., 2019) with VGGish embeddings (Hershey et al., 2017), the Kullback–Leibler divergence (KLD) with PASST audio encoder (Koutini et al., 2021) and CLAP [^2] https://github.com/LAION-AI/CLAP cosine similarity (Elizalde et al., 2023). For music editing evaluations we compute the average L2 distance between the original and edited latent sequences (LPAPS (Iashin & Rahtu, 2021)), FADedit between the distribution of source and edited samples and CLAPedit between the edited audio and the editing prompt. Subjective evaluations relate to (i) overall quality (OVL), and (ii) relevance to the text input (REL), both using a Likert scale (from 1 to 5). Additionally for music editing evaluations we report (iii) editing consistency (CON). Raters were recruited using the Amazon Mechanical Turk platform and all samples were normalized to -14dB LUFS (Series, 2011). For stereo samples objective evaluation the signal is down mixed into mono prior to metrics computation. For subjective ratings we keep the original audio format generated by each model. A screenshot of the evaluation form is presented in appendix A.1.4.
@@ 4.1 Text-guided music editing
We compare MelodyFlow-medium with existing open source music editing implementations, namely MusicGen-melody and AudioLDM 2 with DDPM inversion (following Manor & Michaeli (2024)). The Table 1 presents the main music editing subjective evaluation results. MelodyFlow outperforms both baselines on the quality and text-fidelity axes. MusicGen-melody specifically underperforms consistency-wise while AudioLDM 2 suffers from lower text adherence. Indeed during our listening tests we observe that AudioLDM 2 with DDPM inversion sometimes only generates a distorted version of the original track, hence does not take into account the editing prompt and keeps a strong similarity with the original. This also explains why consistency-wise MelodyFlow lags slightly behind AudioLDM 2. Averaging on the three axes MelodyFlow sets a new baseline for zero-shot music editing at test-time.
4.2 Text-to-music generation
Text-to-music generation performance is reported in the Table 2. For text-to-music qualitative evaluations we compare MelodyFlow to three baselines that also support both generation and editing: MusicGen, AudioLDM 2, Stable-Audio. For MusicGen and AudioLDM 2 we use the available open source implementations and for Stable-Audio we use the public API (as of Wed. May 14 2024, AudioSpark 2.0 model version). MelodyFlow achieves comparable performance with MusicGen, both lagging slightly behind Stable-Audio in terms of human preference. We do not report objective metrics on Stable-Audio as none were reported on the full MusicCaps benchmark Evans et al. (2024a). We do not run any subjective evaluation against MAGNeT but report their objective metrics and latency values. MelodyFlow achieves remarkable efficiency with only 64 inference steps.
[…]
Reference - - - 3.67$\pm$0.10 4.04$\pm$0.10 - -
AudioLDM 2 3.1 $1.20$ $0.31$ 2.79$\pm$0.08 3.40$\pm$0.08 208 18.1
MusicGen-small 3.1 $1.29$ $0.31$ 1500 17.6
MusicGen-medium 3.4 $1.23$ $0.32$ 3.40$\pm$0.08 3.79$\pm$0.07 1500 41.3
Stable-Audio - - - 3.67$\pm$0.08 3.89$\pm$0.07 100 8.0
MAGNeT-small 3.3 $1.12$ $0.31$ 180 4.0
MAGNeT-large 4.0 $1.15$ $0.29$ 180 12.6
MelodyFlow-small 2.8 1.27 0.33 3.27$\pm$0.08 3.83$\pm$0.08 64 1.8
MelodyFlow-medium 3.5 1.30 0.31 3.41$\pm$0.08 3.77$\pm$0.07 64 2.3
4.3 Latent Inversion
4.3.1 Inversion methods
(a) FADedit as a function of $\lambda_{KL}$.
(b) CLAPedit as a function of $\lambda_{KL}$.
(c) LPAPS as a function of $\lambda_{KL}$.
Figure 2: Effect of the regularization weight $\lambda_{KL}$ on the quality (Figure 2(a)) and text-adherence (Figure 2(b)) of music editing. $\epsilon$- and $v$-prediction are compared with or without $c_{orig}$.
We compare MelodyFlow with DDIM and ReNoise in the Figures 2(a), 2(b) and 2(c), as a function of the divergence loss weight $\lambda_{KL}$. During the inversion we use a classifier-free-guidance (CFG) of 0 and employ a CFG of 4 during the regeneration. The choice of zero CFG is meant to prevent divergence during inversion (see A.2.4). For ReNoise and MelodyFlow the predictions are regularized by the weighted KL patch-wise divergence loss $\mathcal{L}_{patchKL}$ of Algorithm 1 and ReNoise additionally uses an autocorrelation loss with $\lambda_{pair}=10$ (Garibi et al., 2024). Both also employ the reversible inversion trajectory estimation while DDIM does not. The Figures show that both MelodyFlow and ReNoise outperform DDIM inversion by a large margin on the three evaluated axes. an optimum can be achieved around $\lambda_{KL}=0.2$ for velocity prediction and around $0.1$ for noise prediction. Overall the quality is better (lower FADedit in the Figure 2(a)) when directly regularizing the velocity prediction. In both cases we observe a higher CLAPedit in the Figure 2(b) when the original text description $c_{orig}$ conditions the inversion process, confirming better text-adherence. This happens at the expense of a higher FADedit compared with unconditional inversion.
4.3.2 Target inversion flow step
In the Figures 3(a), 3(b) and 3(c) we report music editing objective metrics as a function of $T_{edit}$, comparing DDIM inversion with MelodyFlow. The consistency with the source sample is higher (lower LPAPS) with our method than DDIM inversion. The S-shaped FAD curves of the Figure 3(a) indicate an inversion optimum around $T_{edit}=0.06$, correlating with the peak in CLAPedit score.
(a) Editing FADedit vs. $T_{edit}$.
(b) Editing CLAPedit vs. $T_{edit}$.
(c) Editing LPAPS vs. $T_{edit}$.
4.4 Codec bottleneck regularizer
All other things being equal, we ablate on the bottleneck regularizer for a fixed frame rate of 50 Hz by comparing RVQ- (using 4 codebooks of size 2048 each), KL-regularizer (Evans et al., 2024a) and no regularizer at all in the Table 3. Our results indicate optimal reconstruction performance with no regularizer, closely followed by KL. RVQ stands much further away, likely due to the high level of compression enforced by the discretization (despite the significant dictionary size of $2048^{4}=1.7\times 10^{13}$). The same ranking applies for SI-SDR (Le Roux et al., 2019). Regarding text-to-music generation performance, the KL-regularizer outperforms the other options. Overall this shows the KL-regularizer offers the best trade off between reconstruction and generation performance.
Ablating on the codec frame rate with the KL regularizer shows that 5 Hz achieves comparable performance with the 50 Hz codecs trained with RVQ or no regularizer, a 10$\times$ improvement factor. We chose to work with the 20 Hz KL-regularized codec for the 32 kHz mono MelodyFlow-small, as it offers a good trade off between quality and speed. Accounting for the additional information to compress when scaling to 48 kHz stereo, we chose a frame rate of 25 Hz for MelodyFlow-medium.
@@ 4.4 Codec bottleneck regularizer
4.5 FM design
4.5.1 Model training
We compare our FM model design with the baseline implementation of Le et al. (2024), both being trained on the same music latents. The most notable changes are the removal of the infilling objective during training, the change in flow step sampling and the introduction of mini-batch coupling. Table 4 presents the impact of those choices on the last FM model validation MSEloss of the EMA checkpoint, and on the in-domain test FAD (in the text-to-music generation setting). No loss value is reported for the baseline as the infilling objective facilitates the task (hence values are not fairly comparable), and for validation we sample flow steps uniformly regardless of the training sampling scheme. Such infilling objective in Le et al. (2024)’s FM model was designed to handle variable length sequences that are inherent to the speech domain. In our experiments it showed to be detrimental for the model performance, and we know diffusion models can support infilling/outfilling without additional tweaks (Liu et al., 2023a). With all methods combined the in-domain FAD is reduced to 0.39 from 0.53 and consistent with the observed loss decrease, which validates our design.
Table 4: FM model design ablation. FAD (resp. MSE) is reported on the in-domain test (resp. validation) set. Baseline is adapted from (Le et al., 2024) but retrained on our music latents.
Ablation Heads Layers Infill Sampling OT-FM MSEloss $\downarrow$ FADvgg $\downarrow$
baseline 16 24 uniform - .53
-- infilling 16 24 uniform .8596 .50
$+$ sampling 16 24 logit-normal .8484 .44
$+$ batch coupling 16 24 logit-normal .8322 .42
$+$ wider model 18 18 logit-normal .8310 .39
@@ 4.5.1 Model training
4.5.2 Music editing quality/efficiency
We compare DDIM with MelodyFlow’s inversion using a target inversion step of $T_{edit}=0$. FADedit (Figure 4(a)), CLAPedit (Figure 4(b)) and LPAPS (Figure 4(c)) are plotted as a function of the total NFE count (inversion + regeneration included). Quality-wise, the combination of our FM and inversion designs outperform the baseline. Regardless of the FM design choice, DDIM inversion requires as few as 32 NFEs to achieve an acceptable FAD. Our inversion only outperforms after 125 NFEs. On the text-adherence axis, the FM model design alone does not translate in better performance when combined with DDIM inversion. Swapping DDIM with our method shows a different trend, highlighting the benefit of combining FM and inversion methods. Analyzing the consistency with the original sample, again we observe that the regularized inversion plays a more important role than the FM model design: the baseline FM model actually outperforms ours when used in conjunction with DDIM inversion. Overall our method consistently outperform the baseline for 125 NFEs.
@@ 5 Related work
5.1 Audio representation
Recent advancements in neural codecs have seen the application of VQ-VAE on raw waveforms, incorporating a RVQ bottleneck as demonstrated in Zeghidour et al. (2021); Défossez et al. (2022), later refined as per Kumar et al. (2024). Evans et al. (2024a) proposed a modification to this approach by replacing the RVQ with a VAE bottleneck to enhance the modeling of continuous representations. In addition, several recent audio generative models have adopted Mel-Spectrogram latent representations, coupled with a vocoder for reconstruction, as shown in the works of (Ghosal et al., 2023; Liu et al., 2023b; Le et al., 2024).
5.2 Text-to-music generation
Models that operate on discrete representation are presented in the works of (Agostinelli et al., 2023; Copet et al., 2024; Ziv et al., 2023). Agostinelli et al. (2023) proposed a representation of music using multiple streams of tokens, which are modeled by a cascade of transformer decoders conditioned on a joint textual-music representation (Huang et al., 2022b). Copet et al. (2024) introduced a single-stage language model that operates on streams of discrete audio representations, supporting both 32 kHz mono and stereo. Ziv et al. (2023) replaced the language model with a masked generative single-stage non-autoregressive transformer. Schneider et al. (2023); Huang et al. (2023); Liu et al. (2023b) use diffusion models. Schneider et al. (2023) utilized diffusion for both the generation model and the audio representation auto-encoder. Liu et al. (2023b) trained a foundational audio generation model that supports music with latent diffusion, conditioned on autoregressively generated AudioMAE features (Huang et al., 2022a). Evans et al. (2024a, b) proposed an efficient long-form stereo audio generation model based on the latent diffusion of VAE latent representations. This model introduced timing embeddings conditioning to better control the content and length of the generated music.
5.3 Music editing
Lin et al. (2024) proposed a parameter-efficient fine-tuning method for autoregressive language models to support music inpainting tasks. Garcia et al. (2023) developed a masked acoustic modeling approach for music inpainting, outpainting, continuation and vamping. Wu et al. (2023) fine-tuned a diffusion-based music generation model with melody, dynamics and rhythm conditioning. Novack et al. (2024) is a fine-tuning free framework for controlling pre-trained text-to-music diffusion models at inference-time via initial noise latent optimization. Zhang et al. (2024) investigated zero-shot text-guided music editing with conditional latent space and cross attention maps manipulation. Manor & Michaeli (2024) employs DDPM inversion (Huberman-Spiegelglas et al., 2023) for zero-shot unsupervised and text-guided audio editing.